OpenStudy (anonymous):

just to check my math, cos pi/10 = .95 sin pi/10 =.31

OpenStudy (anonymous):

\[[\frac{ 1 }{ 4 }(\cos \frac{ \pi }{ 10 }+i \sin \frac{ \pi }{ 10 })]^{5}\]

OpenStudy (perl):

there is a theorem we can use

OpenStudy (perl):

( r cos theta + i sin theta ) ^n = r^n * cos (theta*n ) + i sin (thetan * n )

OpenStudy (anonymous):

i found something like this, \[\frac{ 1 }{ 4 }^{5}(\cos (5*\frac{ \pi }{10}+i \sin (5*\frac{ \pi }{ 10 })\]

OpenStudy (perl):

yes thats right

OpenStudy (anonymous):

here is my findings, .000976(.000796)+i(.999)

OpenStudy (perl):

what is cos(pi/2) ?

OpenStudy (perl):

cos (5pi/10) = cos(pi/2) = 0

OpenStudy (anonymous):

where did cos(pi/2)come from

OpenStudy (perl):

5pi/10 = pi/2

OpenStudy (anonymous):

so disregard my last equation, \[\frac{ 1 }{ 4 }^{5}(\cos \frac{ \pi }{ 2}+isin \frac{ \pi }{ 2 })\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }^{5}(0+i*1)\]

OpenStudy (anonymous):

final answer would be \[\frac{ 1 }{1024 }i\]